Tennis balls are sold in rectangular box of six balls. Each ball has a radius of 3cm and a volume of 113cubic cm.
a) write down the dimensions of the rectangular box.
b) calculate the volume of the rectangular box
c) calculate the volume of the rectangular box not occupied by the balls
d) calculate the percentage of the volume of the rectangular box which is occupied by the balls. Give your answer to the nearest integer.

Answer by ewatrrr(10682)

 
Hi,
Tennis balls are sold in rectangular box of six balls.
Each ball has a radius of 3cm (d = 6cm) and a volume of 113cubic cm.
a) write down the dimensions of the rectangular box.
6*6cm = 36cm by 6cm by 6cm
b) calculate the volume of the rectangular box = 36cm*6cm*6cm = 1296cm^3
c) calculate the volume of the rectangular box not occupied by the balls
1296cm^3 - 6* 113cm^3 = 618cm^3
d) calculate the percentage of the volume of the rectangular box which
is occupied by the balls. 618/1296 = 48%